A 6 inch personal pizza has 600 calories, with 240 of those from fat. A 12 inch pizza is cut into 8 slices. Estimate the number of calories in one slice of a 12 inch pizza.
Answer #1

Hi

Area of 6 inch pizza/Area of 12 inch pizza = calories of 6 inch pizza/ calories of 12 inch pizza

6^2/12^2 = 600/calories of 12 inch pizza

calories of 12 inch pizza = 600*144/36

= 600*4 = 2400 calories

calories in each slice of pizza = 2400/8

= 300 calories